# Sparse-DeRF: Deblurred Neural Radiance Fields from Sparse View

### v1

@@ IV-B Modulated Gradient Scaling
However, we experimentally found that the limitation is more prominent in sparse view settings since there is less available diversity of viewing direction, which means the projective geometry and epipolar geometry do not properly work for NeRF optimization. Therefore, the gradient scaling seems to be more necessary to the radiance fields from the sparse view setting. Although we tried to apply the technique to the Sparse-DeRF, it does not properly works as demonstrated in the appendix. The reason is that the fixed square function of $J$, which is lower bounded as $1$, does not cover the non-linear parametrized space such as normalized device coordinates (NDC), which is commonly used as well as our work. Another reason is that a strictly fixed shape of the function cannot cover the arrangements of the scene components, which are usually different across the scene even in the same dataset. [51] briefly mentioned the determinant of the jacobian as an additional scaling factor for the value of $J$, but they did not experiment on it. Moreover, even if the additional scaling factor were applied, the shape of the scaling function would remain unchanged and simply in the form of the square function, which does not allow for flexible adaptation to the arrangement of scene components. Therefore, we modulate the shape of the scaling function $J$ to adaptively reflect the scene arrangements and be suitable for NDC.
Our novel gradient scaling function $\hat{J}$ is designed based on three conditions. First, the function should increase from zero at the camera origin, which is a critical condition to avoid the local minima in the initial training phase we mentioned before. Second, the function should be not zero in the far distance, which is set to $1$ in our NDC environment, to ensure the NeRF training. Finally, the function is designed to increase and decrease only once within a given depth range, which makes the function not fluctuate, since it is an intuitively reasonable scenario considering the goal of MGS that alleviates incorrect density mapping in near distance. In addition to the above conditions for proper shape of the scaling function, we further consider designing the function shape when the location of the main objects is focused on the center of the scene and density mapping error that is represented as a painted wall of near or far depth. Following the conditions, the proposed modulated gradient scaling (MGS) function $\hat{J}$ is formulated as
@@ IV-C Perceptual Distillation
However, it is not possible to directly utilize the pre-deblurred images as additional pixel-wise color supervision, due to the lack of 3D consistency. This inconsistency occurs due to the ill-posed property of image deblurring and independent deblurring processing across multi-view images, which generates incoherent deblurred results. In the appendix, we additionally address this issue and present the qualitative comparison of pre-deblurred images and reference images, which are estimated from the DP-NeRF [2] trained with the full view, to reveal the geometric inconsistency issue.
To overcome the intrinsic inconsistency of pre-deblurred images, we address the pre-filtered images as a perceptual prior, which transfers the deblurred texture information by extracting features from rendered patches and deblurred images with a pre-trained feature extractor. Figure 5 simply illustrates the pipeline of our perceptual distillation module.
@@
where $\hat{\varphi}_{ptc}$ and $\bar{\varphi}_{ptc}$ indicate extracted features from each color patches, respectively. Note that, the color patch $\hat{C}_{ptc}$ is rendered by forwarding the NeRF MLPs without blur kernel to perceptually transfer the pre-filtered texture information to the implicit clean radiance fields, which is indicated as patch radiance in Figure 2. Then we apply the perceptual loss $\mathcal{L}_{pd}$ [52] to distill the feature information as
@@ V-B Experimental Sparse View Setting
Scenarios for the Sparse-DeRF are assumed to be three kinds of settings, which are the DeRF from 2-view, 4-view, and 6-view settings unlike existing common sparse-view NeRFs, which usually use 3-view, 6-view, and 9-view settings. The reason is that when using more than 8 images, the joint optimization problem occurs more frequently under 9-view settings as shown in Figure 6. The figure presents the graph of experiments in the Decoration scene where we varied the number of input sparse views from 2-view to 10-view. We also attach the quantitative results in the appendix due to the page limitation. Experimental results reveal that our sparse-view experimental settings are valid since the RegNeRF [6] with DP-kernel shows poor performances under the 9 views.
@@ V-E Evaluation
TABLE I: Average results of novel view synthesis for both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@ V-E1 Quantitative Evaluation
We present the quantitative results of Sparse-DeRF for two different types of blur kernels from Deblur-NeRF [1] and DP-NeRF [2], comparing these results with established baseline methods. The effectiveness of our approach is demonstrated in TABLE I, achieving outstanding performance across entire sparse view settings, regardless of the blur kernel employed. In the Sparse-DeRF results, DN-kernel and DP-kernel denote the blur kernels proposed by the Deblur-NeRF [1] and DP-NeRF [2], respectively. MPR$+$NeRF denotes the naive NeRF model trained solely on color supervision from deblurred images by MPRNet [56], utilizing the reconstruction loss $\mathcal{L}_{recon}$ of Eq.5. The results of MPR$+$NeRF reveal interesting observations and marginal improvements in radiance fields when only employing pre-deblurred images as direct color supervision for training. This tendency emphasizes the 3D inconsistency across the pre-deblurred images, as discussed in Section IV-C. In addition, the poor results of the RegNeRF [6] with and without a blur kernel demonstrate that existing regularization faces difficulty in alleviating the complex joint optimization involving both the blur kernel and radiance fields. We further present an analysis of this optimization issue in Section V-F1 with detailed experimental results. In contrast, Sparse-DeRF demonstrates significant enhancements across all evaluation metrics for both types of blur kernels, indicating its superior ability to represent the DeRF with improved visual quality. In particular, our results exhibit more prominent improvements in real-scene scenarios, although there is non-ideal blur degradation, which occurs due to various real environmental factors. For a comprehensive understanding, we provide an extensive ablation study in Section V-F, incorporating results with both types of blur kernels. Additionally, detailed results for all scenes are appended in the appendix, including synthetic and real scenes.
@@ V-F1 Problem Analysis and Motivation
we present the experimental results to describe that it is difficult to jointly optimize the DeRF and naively apply the previous regularization technique of the NeRF to the DeRF from sparse view setting, utilizing the representative existing regularization method, RegNeRF [6]. We attach the experimental results of the RegNeRF [6] from 2-view, 4-view, and 6-view settings with and without the blur kernel to demonstrate the difficulty of the complex joint optimization problem as we mentioned. TABLE II and Figure 8 present quantitative and qualitative results from 2-view, 4-view, and 6-view, respectively. The blur kernel employed is the rigid blur kernel of DP-NeRF [2], which is described as DP-kernel in the table and figures. To help the reader compare the plausible appearance and dense geometry with the proposed Sparse-DeRF, we attach the rendered color and depth images of our model. The results of the figures demonstrate that the integration of the blur kernel involves a straightforward difficulty in optimizing the high-frequency details and the overall scene geometry simultaneously, as indicated by the visual quality of the rendered color and depth images. Although the presence of a blur kernel enables radiance fields to capture high-frequency details, it causes a geometric distortion with the overall wrong density mapping that resembles the fragmented structure of objects or appearance like the painted wall of near- or far-depth regions. These results and analysis demonstrate that naively applying the regularization of the NeRF from sparse view to the DeRF is not effective. In addition, the images (i) and (ii) in Figure 8 demonstrate that although the overall performance is better without the blur kernels, the model still faces difficulty in modeling high-frequency details. This difficulty makes the necessity of the blur kernel optimization still important, leading to a blurry visual quality across the entire scene. Therefore, motivated by the experimental analysis, we propose Sparse-DeRF, the novel regularization method for optimization of the blur kernel and radiance fields simultaneously. The proposed Sparse-DeRF presents high-quality rendered images with dense geometry and detailed high-frequency texture as shown in Figure 8.
TABLE II: Average results for the complex optimization issue in both synthetic and real scenes from 2-view, 4-view, and 6-view settings. Each color shading represents the best and second results for each experimental setting, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
Blur Kernel  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)  PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$)
RegNeRF [6] $\times$ 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
RegNeRF [6] DP-kernel 13.24 .2162 .6062 12.76 .1836 .6447 16.44 .3657 .4826 13.59 .2221 .5887 21.60 .6162 .3046 18.25 .4439 .4326
Sparse-DeRF (Ours) DP-kernel 15.35 .2904 .5242 15.57 .3114 .5467 21.05 .5776 .2975 20.05 .5178 .3736 24.27 .7255 .2044 22.32 .6283 .2907
@@
TABLE III: Ablation experimental results from 2-view, 4-view, and 6-view according to the proposed geometric constraints and perceptual prior, denoted in the table as surface smoothness (SS), modulated gradient scaling (MGS), and perceptual distillation (PD), respectively. We present separate results for our model with both types of kernels, DP-kernel [2] and DN-kernel [1]. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@ V-F2 Ablation of Each Component
We demonstrate the effectiveness of the Sparse-DeRF’s each component through comprehensive ablation studies, presenting both quantitative and qualitative results. In TABLE III and Figure 9, we present independent quantitative and qualitative results from SS, MGS, and PD, where each component is individually applied to observe the influence of each method. The results include several combinations of the components to reveal the complement effect of each proposed method. The experiments are conducted for entire 2-view, 4-view, and 6-view settings, providing a different performance depending on two types of blur kernels, DN-kernel [1] and DP-kernel [2]. Our model shows superior results in predicting both 3D geometric and appearance details precisely, demonstrating the enhanced evaluation results.
@@ V-F3 Ablation Analysis
Quantitative ablation results demonstrate that our model with full components shows the best results in the real-scene dataset. The results reveal that MGS plays an important role across the entire 2-view, 4-view, and 6-view settings among geometric constraints although each constraint enhances the geometric accuracy. In Figure 9, the importance of MGS is more clearly prominent with qualitative results as color and depth images. If MGS is not applied, we can see that the geometry of the scene is not accurately captured or there are many density artifacts. On the other hand, PD seems to be not effective without geometric constraints in 2-view and 4-view settings as we can figure out in TABLE III and Figure 9. The reason is that the NeRF has more difficulty in predicting the correct geometry with only pre-deblurred images due to the inherent 3D inconsistency, which makes perceptual prior not effective. These difficulties become more severe as the number of views decreases. We can demonstrate that a certain level of accurate geometry should be achieved before applying perceptual distillation.
TABLE IV: Average quantitative comparison results between naive gradient scaling of [51] and our MGS in both synthetic and real scenes from 2-view, 4-view, and 6-view settings. The two types of kernels we utilize in this paper are indicated as DP-kernel and DN-kernel, the kernels of DP-NeRF [2] and Deblur-NeRF [1], respectively. Color shading represents the better result.
@@ V-F4 Comparison to Naive Gradient Scaling
We present experimental results of quantitative and qualitative comparison between our proposed MGS and naive gradient scaling of [51] from 2-view, 4-view, and 6-view settings in TABLE IV and Figure 10. We compare the effectiveness of our MGS for two types of kernels we utilize in our paper, which are the kernels of Deblur-NeRF [1] and DP-NeRF [2]. The results describe our MGS outperforms the naive gradient scaling in terms of both quantitative and qualitative performance across the entire experimental setting. Specifically, depth images demonstrate that our MGS more effectively helps the model to predict the accurate geometry than naive gradient scaling. Quantitative results for the entire scene are attached in the appendix.
V-F5 Inconsistency of Pre-deblurred Images

### v2

@@ IV-B Modulated Gradient Scaling
However, we experimentally found that the limitation is more prominent in sparse view settings since there is less available diversity of viewing direction, which means the projective geometry and epipolar geometry do not properly work for NeRF optimization. Therefore, the gradient scaling seems to be more necessary to the radiance fields from the sparse view setting. In addition, we believe the proper gradient scaling can be an implicit geometric guidance in sparse view setting. Although we tried to apply the technique to the Sparse-DeRF, it does not properly works as demonstrated in the appendix. The reason is that the fixed square function of $J$, which is lower bounded as $1$, does not cover the non-linear parametrized space such as normalized device coordinates (NDC), which is commonly used as well as our work. Another reason is that a strictly fixed shape of the function cannot cover the arrangements of the scene components, which are usually different across the scene even in the same dataset. [60] briefly mentioned the determinant of the jacobian as an additional scaling factor for the value of $J$, but they did not experiment on it. Moreover, even if the additional scaling factor were applied, the shape of the scaling function would remain unchanged and simply in the form of the square function, which does not allow for flexible adaptation to the arrangement of scene components or guide the rough geometry of the radiance fields. Therefore, we modulate the shape of the scaling function $J$ to adaptively reflect the scene arrangements and guide the rough geometry while suitable for NDC, since we use NDC for our experiments.
Our novel gradient scaling function $\hat{J}$ is designed based on three conditions. First, the function should increase from zero at the camera origin, which is a critical condition to avoid the local minima in the initial training phase we mentioned before. Second, the function should not be zero in the far distance, which is set to $1$ in our NDC environment, to ensure the NeRF training. Second, the function should not be zero in the far distance, which is set to $1$, to ensure the NeRF training. Finally, the function is designed to increase and decrease only once within a given depth range, which makes the function not fluctuate, since it is an intuitively reasonable scenario considering the goal of MGS that alleviates incorrect density mapping in near distance. In addition to the above conditions for proper shape of the scaling function, we further consider designing the function shape when the location of the main objects is focused on the center of the scene and density mapping error that is represented as a painted wall of near or far depth. Following the conditions, the proposed modulated gradient scaling (MGS) function $\hat{J}$ is formulated as
@@ IV-C Perceptual Distillation
However, it is not possible to directly utilize the pre-deblurred images as additional pixel-wise color supervision, due to the lack of 3D consistency. This inconsistency occurs due to the ill-posed property of image deblurring and independent deblurring processing across multi-view images, which generates incoherent deblurred results. In Section V-F6, we additionally address this issue and present the qualitative comparison of pre-deblurred images and reference images, which are estimated from the DP-NeRF [2] trained with the full view, to reveal the geometric inconsistency issue.
To overcome the intrinsic geometric inconsistency of pre-deblurred images, we address the pre-deblurred images as a perceptual prior, which transfers the only deblurred texture information by extracting features from rendered patches and deblurred images with a pre-trained feature extractor. Figure 5 simply illustrates the pipeline of our perceptual distillation module. To the best of our knowledge, utilizing pre-deblurred images as an additional prior in deblurring radiance fields is a first attempt.
@@
where $\hat{\varphi}_{ptc}$ and $\bar{\varphi}_{ptc}$ indicate extracted features from each color patches, respectively. Note that, the color patch $\hat{C}_{ptc}$ is rendered by forwarding the NeRF MLPs without blur kernel to perceptually transfer the pre-deblurred texture information to the implicit clean radiance fields, which is indicated as patch radiance in Figure 2. Then we apply the perceptual loss $\mathcal{L}_{pd}$ [61] to distill the feature information as
@@ V-B Experimental Sparse View Setting
Scenarios for the Sparse-DeRF are assumed to be three kinds of settings, which are the DeRF from 2-view, 4-view, and 6-view settings unlike existing common sparse-view NeRFs, which usually use 3-view, 6-view, and 9-view settings. The reason is that the joint optimization problem occurs more frequently under 9-view settings as shown in Figure 6. The figure presents the graph of experiments in the Decoration scene where we varied the number of input sparse views from 2-view to 10-view. We also attach the quantitative results in the appendix. Experimental results reveal that our sparse-view experimental settings are valid since the RegNeRF [6] with DP-kernel shows poor performances under the 9 views.
@@ V-E Evaluation
TABLE I: Average results on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@ V-E1 Quantitative Evaluation
We present the quantitative results of Sparse-DeRF for two different types of blur kernels from Deblur-NeRF [1] and DP-NeRF [2], comparing these results with established baseline methods. The effectiveness of our approach is demonstrated in TABLE I, achieving outstanding performance across entire sparse view settings, regardless of the blur kernel employed. In the Sparse-DeRF results, DN-kernel and DP-kernel denote the blur kernels proposed by the Deblur-NeRF [1] and DP-NeRF [2], respectively. NeRF$+$MPR denotes the naive NeRF model trained solely on color supervision from deblurred images by MPRNet [65], utilizing the reconstruction loss $\mathcal{L}_{recon}$ of Eq.5, which is samely utilized in RegNeRF+MPR experiment. The results of the experiment with MPR [65] reveal interesting observations and marginal improvements in radiance fields when only employing pre-deblurred images as direct color supervision for training. This tendency even prominent when observing the results of applying MPR [65] to both NeRF (2 row in TABLE I) and RegNeRF (6 row in TABLE I). Regardless of the presence or absence of a regularizer, the improvement is marginal and the performance of perceptual quality(LPIPS score) even decreases when a larger number of images is used due to the inconsistency as we mentioned above. In the RegNeRF+MPR experiments, we observe a slight increase in PSNR in the 2-view setting. In this case, we expect the somewhat cleaner image from the pre-deblurred image appears to be the advantage, since there are significant lack of information in the experimental environment. Nonetheless, perceptual quality significantly deteriorates with poor LPIPS scores, which demonstrates the pre-deblurred images are not quite helpful to enhance the quality of the reconstructed radiance fields. This tendency emphasizes the 3D inconsistency across the pre-deblurred images, as discussed in Section IV-C.
In addition, the poor results of the RegNeRF [6] with and without a blur kernel demonstrate that existing regularization faces difficulty in alleviating the complex joint optimization involving both the blur kernel and radiance fields. We further present an analysis of this optimization issue in Section V-F1 with detailed experimental results. In contrast, Sparse-DeRF demonstrates significant enhancements across all evaluation metrics for both types of blur kernels, indicating its superior ability to represent the DeRF with improved visual quality. In particular, our results exhibit more prominent improvements in real-scene scenarios, although there is non-ideal blur degradation, which occurs due to various real environmental factors. For a comprehensive understanding, we provide an extensive ablation study in Section V-F, incorporating results with both types of blur kernels. Additionally, detailed results for all scenes are appended in the appendix, including synthetic and real scenes.
@@ V-F1 Problem Analysis and Motivation
we present the experimental results to describe that it is difficult to jointly optimize the DeRF and naively apply the previous regularization technique of the NeRF to the DeRF from sparse view setting, utilizing the representative existing regularization method, RegNeRF [6]. We attach the experimental results of the RegNeRF [6] from 2-view, 4-view, and 6-view settings with and without the blur kernel to demonstrate the difficulty of the complex joint optimization problem as we mentioned. The results of the RegNeRF w/ and w/o DP-kernel in TABLE I (5, 6, 8 row) and Figure 8 present quantitative and qualitative evidence of complex optimization from sparse view setting and shows the outstanding performance of our model. To help the reader compare the plausible appearance and dense geometry with the proposed Sparse-DeRF, we attach the rendered color and depth images of our model. The results of the figures demonstrate that the integration of the blur kernel involves a straightforward difficulty in optimizing the high-frequency details and the overall scene geometry simultaneously, as indicated by the visual quality of the rendered color and depth images. Although the presence of a blur kernel enables radiance fields to capture high-frequency details, it causes a geometric distortion with the overall wrong density mapping that resembles the fragmented structure of objects or appearance like the painted wall of near- or far-depth regions. These results and analysis demonstrate that naively applying the regularization of the NeRF from sparse view to the DeRF is not effective. In addition, the images (i) and (ii) in Figure 8 demonstrate that although the overall performance is better without the blur kernels, the model still faces difficulty in modeling high-frequency details. This difficulty makes the necessity of the blur kernel optimization still important, leading to a blurry visual quality across the entire scene. Therefore, motivated by the experimental analysis, we propose Sparse-DeRF, the novel regularization method for optimization of the blur kernel and radiance fields simultaneously. The proposed Sparse-DeRF presents high-quality rendered images with dense geometry and detailed high-frequency texture as shown in Figure 8.
@@
TABLE II: Average results of the ablation study for the comparison of the proposed geometric constraints and perceptual prior, denoted in the table as surface smoothness (SS), modulated gradient scaling (MGS), and perceptual distillation (PD). The experiement is conducted on novel view synthesis for both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best, and third best result for each experimental setting, respectively.
@@ V-F2 Ablation of Each Component
We demonstrate the effectiveness of the Sparse-DeRF’s each component through comprehensive ablation studies, presenting both quantitative and qualitative results. In TABLE II and Figure 9, we present independent quantitative and qualitative results from SS, MGS, and PD, where each component is individually applied to observe the influence of each method. The results include several combinations of the components to reveal the complement effect of each proposed method. The experiments are conducted for entire 2-view, 4-view, and 6-view settings, providing a different performance depending on two types of blur kernels, DN-kernel [1] and DP-kernel [2]. Our model shows superior results in predicting both 3D geometric and appearance details precisely, demonstrating the enhanced evaluation results.
@@ V-F3 Ablation Analysis
Quantitative ablation results demonstrate that our model with full components shows the best results in the real-scene dataset. The results reveal that MGS plays an important role across the entire 2-view, 4-view, and 6-view settings among geometric constraints although each constraint enhances the geometric accuracy. In Figure 9, the importance of MGS is more clearly prominent with qualitative results as color and depth images. If MGS is not applied, we can see that the geometry of the scene is not accurately captured or there are many density artifacts. On the other hand, PD seems to be not effective without geometric constraints in 2-view and 4-view settings as we can figure out in TABLE II and Figure 9. The reason is that the NeRF has more difficulty in predicting the correct geometry with only pre-deblurred images due to the inherent 3D inconsistency, which makes perceptual prior not effective. These difficulties become more severe as the number of views decreases. We can demonstrate that a certain level of accurate geometry should be achieved before applying perceptual distillation.
TABLE III: Average results between naive gradient scaling of [60] and our MGS. The experiements are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Color shading represents the better result.
@@
TABLE IV: Average results of the comparison with the model that uses depth-prior. The experiments are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second best results.
[ 2-view ] [ 4-view ] [ 6-view ]
SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.08 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
Figure 11: Qualitative results on Puppet, Ball, and Basket scene from 2-view, 4-view, and 6-view settings with the SparseNeRF [52] that use depth-prior. Even with inaccurate depth priors, it can be helpful in capturing rough geometry in extreme scenarios like a 2-view setting. However, when using 4 or 6 images, the multi-view inconsistency of the predicted depth prior leads to rather inaccurate geometry predictions, especially in thin or detailed structures. Note that, reference image is rendered from DP-NeRF [2] trained with full-view images for comparison.
TABLE V: Average results of the comparison depending on usage of the pre-deblurred images on various models. The experiements are conducted on both synthetic and real scenes obtained from 2-view, 4-view, and 6-view settings. Each color shading represents the best, second, and third result for each experimental setting, respectively.
[ 2-view ] [ 4-view ] [ 6-view ]
NeRF + MPR [65] 15.16 .3006 .5595 14.38 .2594 .6019 20.00 .5381 .3956 18.89 .4829 .4484 21.72 .5999 .3629 20.60 .5513 .4010
RegNeRF [6] 14.60 .2849 .5869 15.49 .2997 .5888 18.39 .4600 .4704 18.44 .4326 .4852 19.69 .5249 .4165 19.65 .4790 .4583
RegNeRF [6] + MPR [65] 14.99 .2980 .5690 15.84 .3040 .5818 18.54 .4678 .4593 18.66 .4425 .4804 19.89 .5339 .4071 19.64 .4753 .4597
SparseNeRF [52] 18.12 .4729 .4884 18.01 .4683 .5245 21.08 .6105 .4042 20.69 .5909 .4450 21.77 .6369 .3903 21.28 .6189 .4250
SparseNeRF [52] + MPR [65] 18.14 .4850 .4531 17.65 .4501 .5126 21.67 .6623 .3381 20.88 .6165 .4132 22.47 .6995 .3156 21.53 .6479 .3935
Figure 12: Qualitative results on Heron scene from 4-view setting for the comparison of NeRF [19], RegNeRF [6] the SparseNeRF [52], and ours depending on the pre-deblurred images by using MPRNet [65]. Note that, reference image is rendered from DP-NeRF [2] trained with full-view images for comparison.
@@ V-F4 Comparison to Naive Gradient Scaling
We present experimental results of quantitative and qualitative comparison between our proposed MGS and naive gradient scaling of [60] from 2-view, 4-view, and 6-view settings in TABLE III and Figure 10. We compare the effectiveness of our MGS for two types of kernels we utilize in our paper, which are the kernels of Deblur-NeRF [1] and DP-NeRF [2]. The results describe our MGS outperforms the naive gradient scaling in terms of both quantitative and qualitative performance across the entire experimental setting. Specifically, depth images demonstrate that our MGS more effectively helps the model to predict the accurate geometry than naive gradient scaling. Quantitative results for the entire scene are attached in the appendix. Moreover, since experiments on the DeblurNeRF [1] dataset were conducted in NDC by default, we adopted hyperparameters of MGS to be optimized for NDC. However, since MGS provides geometric guidance for the scene, it naturally works in other coordinate systems as well. To validate this, we also included experimental results in the appendix demonstrating its effectiveness in the world coordinate system.
V-F5 Challenges in Utilizing Depth Priors
We further address this inconsistency issue in relation to depth priors. The different blur in each image lead to accurate and multi-view inconsistent depth estimation, which is not proper to be used as additional prior for predict radiance fields with clean geometric structure. In TABLE IV and Figure 11, we present the quantitative and qualitative results of the proposed Sparse-DeRF and SparseNeRF [52], which is the representative depth-prior based model. The results show that as the number of images increases, the improvement diminishes and scene geometry is distorted, revealing the issue of inconsistent depth priors. In extreme cases like a 2-view setting, where it is challenging to obtain even rough geometry information, even inaccurate depth maps can be somewhat helpful. In addition, since only two images are used, the inconsistency issue is less pronounced, resulting in relatively better structured geometry. However, in 4- or 6-view settings, relying on inconsistent depth prior can make the prediction of a clean and accurate scene geometry even more difficult, particularly for thin or detailed structures as we can observe in the emphasized region in Figure 11. In particular, the qualitative experimental results with 4 views further support the inconsistency issue, as the significant difference in perceptual quality, reflected by LPIPS, provides additional evidence. It is once again clear that qualitative evaluation and LPIPS performance are crucial in the deblurring task, while PSNR and SSIM are important, as mentioned in DP-NeRF [2]. Furthermore, in the appendix, we included the experimental results of SparseNeRF [52] with depth maps predicted by Depth-Anything V2 [68], one of the state-of-the-art monocular depth estimation models, which also exhibits the experimental tendency and multi-view inconsistency as well. Therefore, to design a general sparse view model that works well across 2-, 4-, and 6-views, we adopted an experimental setup that does not utilize depth priors. Quantitative results for the entire scene are attached in the appendix.
V-F6 Inconsistency of Pre-deblurred Images